) Find the real part of the following expression: (1 + i) e(1-i)t + (1 - i)e(1+i)t

Validate with Matlab
Answer #1

Solution :- To find Real part of this Expression (1+i) e l-i) + (1-1) e11+)+. = (1+i) et eitt (1-1) et eit = (1+i) et 163 + -